Peter has $20 and Paul has $212. when they each given $x,Paul has 5 times as much as Peter. Write down an equation in x and solve it.?
Answer by FrankM(1040)

5(20+x)=212+x
100+5x=212+x
100+4x=212
4x=212-100
4x=112
x=28
so ....
20+28= 48 and 212+28=240 240 is 5 times 48